# High-contrast double Bragg interferometry via detuning control

### v1

@@ II.2.4 Mach-Zehnder protocol with optimal control theory
The final detuning-controlled Mach-Zehnder strategy focuses on optimizing the mirror pulse—identified as the performance bottleneck in the DS-DBD protocol—using optimal control theory (OCT), using QCTRL’s Boulder Opal package Ball et al. (2021). Unlike SBD mirror pulse, the DBD mirror pulse involves a four-photon transition $|\pm 2\hbar k_{L}\rangle\rightarrow|0\hbar k_{L}\rangle\rightarrow|\mp 2\hbar k_{L}\rangle$ making it more sensitive to finite momentum spread and nonzero COM momentum Szigeti et al. (2012); Müller et al. (2008); Chiow et al. (2011); Stolzenberg et al. (2025). As illustrated in Fig. 1, a C-DBD mirror selectively filters and only reflects the central momentum component. This sensitivity is mitigated by jointly optimizing a smooth time-dependent detuning $\Delta(t)$ and all three Gaussian pulse parameters $(\Omega_{M},\,\tau_{M},\,t_{0})$, yielding an OCT mirror pulse with $(\Omega_{M},\,\tau_{M},\,t_{0})=(2.502\,\omega_{\mathrm{rec}},\,1.829\,\omega_{\mathrm{rec}}^{-1},\,3.879\,\omega_{\mathrm{rec}}^{-1})$ and an efficiency of $\eta_{M}=99.806\%$ (see Supplementary Material for detuning details), representing a substantial improvement over the DS mirror pulse (see Fig. 3). The resulting hybrid “OCT” sequence combines this mirror with two DS beam-splitters, separated by interrogation time $T$.
@@ III Contrast of double Bragg Mach-Zehnder interferometers
For non-ideal beam-splitter or mirror operations, the contrast is generally less than unity, and the output signal typically contains multiple Fourier components Kirsten-Siemß et al. (2023). However, the contrast can still be defined, analogous to the ideal case, as the population difference between the first two non-trivial extrema of the output signal when scanning the interrogation time $T$:
@@
where $T_{max}$ and $T_{min}$ denote the first non-trivial maximum and minimum of $P_{\pm 2\hbar k}(g,T)\equiv P_{2}(g,T)+P_{3}(g,T)$ at a fixed $g$ as illustrated by the contrast extraction in Fig. 5(b–c). For instance, Fig. 5(b) shows that the contrast extracted for the DS-DBD strategy for an initial input state with a momentum width of $\sigma_{p}=0.05\hbar k_{L}$ and COM momentum $p_{0}=0$ without polarization error is $\mathbf{C}=97\%$. Fig. 5(c) shows the detrimental effect of a large initial COM momentum, $p_{0}=0.1\hbar k_{L}$, on the contrast, where the contrast is reduced to $\mathbf{C}=83\%$ for a narrower momentum width of $\sigma_{p}=0.01\hbar k_{L}$ without polarization error. In both cases, the results of the five-level S-matrix theory agrees perfectly with the exact numerical solutions of the Schrödinger equation based on the second-order Suzuki-Trotter decomposition Fitzek et al. (2020); Suzuki (1990) with a momentum truncation up to $\pm 14.5\hbar k_{L}$. The single Fourier fringe model (Eq. (9)) applies only in the ideal regime of small initial COM momentum and momentum width, where pulse efficiency remains nearly unaffected.
@@
We now turn to the main results, comparing the contrast robustness of different detuning-control strategies against three key experimental imperfections: atomic momentum width $\sigma_{p}$ [Fig. 5(d)], initial center-of-mass momentum $p_{0}$ [Fig. 5(e)], and polarization error $\varepsilon_{pol}$ [Fig. 5(f)]. Under typical experimental conditions with well-controlled polarization errors ($\varepsilon_{pol}<3\%$), we find the performance ranking to be: $\text{OCT}>\text{DS-DBD}>\text{C-DBD}\approx\text{CD-DBD}$. The relative improvements of the DS-DBD and OCT protocols over C-DBD (or CD-DBD) depend on the atomic momentum width (set by the 1D effective temperature), the COM momentum (influenced by laser alignment and pulse timing), and the polarization properties of the optical beams. For a typical atom-chip-generated BEC with a 1D effective temperature of $2\,\text{nK}$ Deppner et al. (2021), corresponding to a momentum width of approximately $\sigma_{p}=0.10\hbar k_{L}$ (for $^{87}\text{Rb}$ atoms at a wavelength of $780\,\text{nm}$), the DS-DBD and OCT strategies yield relative contrast improvements of $3.4\%$ and $12.0\%$, respectively, over the C-DBD protocol (assuming perfect polarization and zero COM momentum). For a BEC after evaporative cooling with a 1D temperature below $500\,\text{pK}$ Leanhardt et al. (2003), or a momentum width about $0.05\hbar k_{L}$, the improvements reduce to $1.3\%$ (for DS-DBD) and $3.8\%$ (for OCT). In the state-of-the-art case of delta-kick collimated BECs with an effective temperature below $40\,\text{pK}$ Deppner et al. (2021); Kovachy et al. (2015), corresponding to $\sigma_{p}=0.014\hbar k_{L}$, the improvements are $0.4\%$ and $0.5\%$, respectively. With well-controlled polarization error and initial COM momentum, the DS-DBD strategy maintains contrast above $90\%$ up to $\sigma_{p}\leq 0.097\,\hbar k_{L}$, while the OCT strategy remains above $95\%$ contrast up to $\sigma_{p}\leq 0.132\,\hbar k_{L}$.
@@
Figure 6: Contrast as a function of relative peak lattice depth fluctuations for different detuning-control strategies. Symbols with error bars represent exact numerical simulations; shaded regions denote theoretical confidence intervals from the five-level S-matrix model.
@@ IV Conclusion
In summary, we have proposed and analyzed high-contrast Mach-Zehnder atom interferometers based on double Bragg diffraction under external acceleration. We introduced a tri-frequency laser scheme to compensate for differential Doppler shifts, enabling efficient beam-splitting and mirror operations. To mitigate contrast loss from experimental imperfections, such as momentum spread, center-of-mass motion, polarization errors, and peak lattice depth fluctuations, we compared the robustness of four optimized detuning-control strategies: C-DBD, CD-DBD, DS-DBD and OCT. We found that the DS-DBD strategy maintains contrast above $90\%$ for atom-chip-generated Bose-Einstein condensates, while the OCT strategy achieves the highest contrast above $95\%$ under all realistic experimental conditions. We also developed a five-level S-matrix model that accurately reproduces the interferometer output signals compared to exact numerical solutions, accounting for imperfect beam-splitter and mirror pulses as well as diffraction phases due to parasitic paths. These strategies are easily adaptable to varying experimental constraints, and alternative detection schemes—such as near-field or mid-field readout—can be naturally incorporated into the five-level S-matrix model via restricted summation indices in Eq. (6). Our results provide practical methods for enhancing contrast in DBD-based interferometers, bringing their performance on par with two-level Raman schemes and enabling their applications in high-precision quantum sensing and fundamental physics tests.
@@
The online Supplementary Material contains two sections. Sec. S1 provides a detailed derivation of the unitary transformation from the laboratory-frame DBD Hamiltonian (Eq. (S1)) to the twin-lattice center-of-mass frame (Eq. (3)). Sec. S2 defines the beam-splitter and mirror pulse efficiencies, evaluates their robustness under experimental imperfections, and outlines the cost function used in the OCT optimization.
@@ acknowledgement
Supplementary Material: High-contrast double Bragg interferometry via detuning control
S1 Tri-frequency DBD Hamiltonian in the twin-lattice Center-of-mass frame
Here, we provide detailed derivation of the Hamiltonian for the tri-frequency double Bragg gravimeter, as presented in Eq. (3) of the main text. We begin with the lab-frame Hamiltonian, copied from Eq. (1) of the paper:
$\Big{\{}\cos\big{[}\Delta\omega(t)t\big{]}+\varepsilon_{pol}\Big{\}}-mg\hat{z}$, (S1)
where $\nu_{D}(t)=2k_{L}a_{L}t$ is the linear Doppler detuning. For brevity, we define $C(t,\varepsilon_{pol})\equiv\cos\big{[}\Delta\omega(t)t\big{]}+\varepsilon_{pol}$ for use below. We claim that this Hamiltonian can be transformed into the center-of-mass (COM) frame of the twin Bragg lattices using the following time-dependent unitary transformation:
$U=e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}-i\frac{\hat{z}}{\hbar}ma_{L}t+i\frac{\Phi(t)}{\hbar}},$ (S2)
with a properly chosen phase $\dot{\Phi}(t)=m(\frac{1}{2}a_{L}-g)\frac{1}{2}a_{L}t^{2}$ and a constant twin-lattice acceleration $a_{L}$.
To show this, let us denote the transformed Hamiltonian in the Bragg twin-lattice COM frame as $H_{COM}(t)$, which is given by
$H_{COM}(t)$ $=UH_{lab}(t)U^{\dagger}+i\hbar\dot{U}U^{\dagger}.$ (S3)
Since the exponent of $U$ contains non-commutative observables $\hat{z}$ and $\hat{p}$ with their commutator given by $\hat{z},\,\hat{p}]=i\hbar$, we apply the Baker–Campbell–Hausdorff formula Campbell (1896, 1897); Rossmann (2002) to calculate both the adjoint action of $U$ on $H_{lab}(t)$ and the time derivative of $U$. We calculate the two parts separately:
$UH_{lab}(t)U^{\dagger}$
$=$ $\frac{1}{2m}U\hat{p}^{2}U^{\dagger}+2\hbar\Omega(t)C(t,\varepsilon_{pol})\times$
$U\cos\Big{(}2k_{L}\hat{z}-\int\!\nu_{D}(t)\,dt\Big{)}U^{\dagger}-mgU\hat{z}U^{\dagger}$
$=$ $\frac{1}{2m}\big{(}U\hat{p}U^{\dagger}\big{)}^{2}+2\hbar\Omega(t)C(t,\varepsilon_{pol})\times$
$\cos\Big{(}2k_{L}\big{(}U\hat{z}U^{\dagger}-\frac{1}{2}a_{L}t^{2}\big{)}\Big{)}-mgU\hat{z}U^{\dagger}$
[…]
$-mg\Big{(}\hat{z}+\frac{1}{2}a_{L}t^{2}\Big{)},$ (S4)
where we have used the relations $U\hat{p}U^{\dagger}=\hat{p}+ma_{L}t$ and $U\hat{z}U^{\dagger}=\hat{z}+\frac{1}{2}a_{L}t^{2}$ in the last step. The second part involving the time derivative of $U$ is given by
$i\hbar\dot{U}U^{\dagger}$
$=$ $i\hbar\frac{d}{dt}\Big{(}e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}-i\frac{\hat{z}}{\hbar}ma_{L}t+i\frac{\Phi(t)}{\hbar}}\Big{)}U^{\dagger}$ (S5)
$=$ $i\hbar\frac{d}{dt}\Big{(}e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}}e^{-i\frac{\hat{z}}{\hbar}ma_{L}t}e^{i\frac{m}{4\hbar}a_{L}^{2}t^{3}+i\frac{\Phi(t)}{\hbar}}\Big{)}U^{\dagger}$ (S6)
$=$ $-a_{L}t\hat{p}+e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}}ma_{L}\hat{z}e^{-i\frac{\hat{z}}{\hbar}ma_{L}t}e^{i\frac{m}{4\hbar}a_{L}^{2}t^{3}+i\frac{\Phi(t)}{\hbar}}U^{\dagger}$
$-\frac{3}{4}ma_{L}^{2}t^{2}-\dot{\Phi}(t)$
$=$ $-a_{L}t\hat{p}+ma_{L}\hat{z}+ma_{L}\Big{[}e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}},\hat{z}\Big{]}\times$
$e^{-i\frac{\hat{z}}{\hbar}ma_{L}t}e^{i\frac{m}{4\hbar}a_{L}^{2}t^{3}+i\frac{\Phi(t)}{\hbar}}U^{\dagger}-\frac{3}{4}ma_{L}^{2}t^{2}-\dot{\Phi}(t)$ (S7)
$=$ $-a_{L}t\hat{p}+ma_{L}\hat{z}-\frac{1}{2}ma_{L}^{2}t^{2}+\frac{1}{2}mga_{L}t^{2},$ (S8)
where the relation $e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}-i\frac{\hat{z}}{\hbar}ma_{L}t}=e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}}e^{-i\frac{\hat{z}}{\hbar}ma_{L}t}e^{i\frac{m}{4\hbar}a_{L}^{2}t^{3}}$ is applied from line (S5) to line (S6) and the relation $e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}},\hat{z}]=\frac{1}{2}a_{L}t^{2}e^{i\frac{\hat{p}}{2\hbar}a_{L}t^{2}}$ is used from line (S7) to line (S8). Combining the two parts together, one recovers Eq. (3) in the paper:
$H_{COM}(t)=$ $\frac{\hat{p}^{2}}{2m}+2\hbar\Omega(t)\cos(2k_{L}\hat{z})C(t,\varepsilon_{pol})$
$-m(g-a_{L})\hat{z}$, (S9)
with $C(t,\varepsilon_{pol})=\cos\big{[}\Delta\omega(t)t\big{]}+\varepsilon_{pol}$.
S2 Efficiencies of beam-splitter and mirror pulses for different detuning-control strategies
Here, we provide additional details regarding the beam-splitter (BS) and mirror (M) pulse efficiencies introduced in the main text. For a given quasi-momentum $p\in[-\hbar k_{L},\hbar k_{L}]$ within the first Brillouin zone, the efficiency of a double Bragg beam-splitter pulse, $\mathcal{F}_{BS}(p)$, is defined as
$\mathcal{F}_{BS}(p)=P_{|p\rangle\rightarrow|p+2\hbar k_{L}\rangle}+P_{|p\rangle\rightarrow|p-2\hbar k_{L}\rangle},$ (S10)
where $P_{|p\rangle\rightarrow|p\pm 2\hbar k_{L}\rangle}$ denotes the transition probability from the initial state $|p\rangle$ to the final state $|p\pm 2\hbar k_{L}\rangle$, respectively. Similarly, the right and left efficiencies of a double Bragg mirror pulse, $\mathcal{F}_{M}^{\pm}(p)$, are defined as
$\mathcal{F}_{M}^{+}(p)$ $=P_{|p+2\hbar k_{L}\rangle\rightarrow|p-2\hbar k_{L}\rangle},$ (S11)
$\mathcal{F}_{M}^{-}(p)$ $=P_{|p-2\hbar k_{L}\rangle\rightarrow|p+2\hbar k_{L}\rangle}.$ (S12)
Due to the parity symmetry of the double Bragg Hamiltonian in the absence of gravity Li et al. (2024), the left and right mirror efficiencies are related by $\mathcal{F}_{M}^{-}(p)=\mathcal{F}_{M}^{+}(-p)$. Without loss of generality, we choose to characterize mirror performance using the right-going efficiency $\mathcal{F}_{M}^{+}(p)$. For an initial state $|\psi(t=0)\rangle=\int\!dp\,\psi(p)|p\rangle$ with compact support in the first Brillouin zone, the integrated beam-splitter and mirror efficiencies are defined as:
$\eta_{BS}$ $=\int_{-\hbar k_{L}}^{\hbar k_{L}}\!\!\!dp|\psi(p)|^{2}\mathcal{F}_{BS}(p),$ (S13)
$\eta_{M}$ $=\int_{-\hbar k_{L}}^{\hbar k_{L}}\!\!\!dp|\psi(p)|^{2}\mathcal{F}_{M}^{+}(p).$ (S14)
S2.1 C-DBD beam-splitter and mirror pulses
The conventional double Bragg (C-DBD) beam-splitter pulse is a temporal Gaussian with parameters $(\Omega_{BS},\tau_{BS},t_{0})=(2.0\,\omega_{\mathrm{rec}},0.47\,\omega_{\mathrm{rec}}^{-1},\,0)$ where $t_{0}$ stands for center of the pulse (fixing the carrier-envelop phase at the pulse center), and a detuning $\Delta=0$. The 2-dimensional (2D) landscape of the conventional double Bragg beam-splitter efficiency versus quasi-momentum and polarization error, i.e., $\mathcal{F}_{BS}(p,\varepsilon_{pol})$, is shown in Fig. S1(a).
Figure S1: Performance of the C-DBD beam-splitter pulse. (a) Landscape of the C-DBD beam-splitter efficiency $\mathcal{F}_{BS}(p,\varepsilon_{pol})$. (b) Transition probabilities to different momentum states for the C-DBD beam-splitter pulse with an input momentum state $|p\rangle$ with no polarization error $\varepsilon_{pol}=0$.
In Fig. S1(b), we plot the transition probabilities to different momentum states with an input momentum state $|p\rangle$ versus quasi-momentum $p$ for a fixed polarization error $\varepsilon_{pol}=0$ predicted by the five-level theory (5-LS) developed in Ref. Li et al. (2024), and compare it with the exact numerical solutions in position space using a very narrow momentum distribution $\sigma_{p}=0.01\hbar k_{L}$ centered around $p_{0}=0$ to approximate the momentum eigenstate $|p\rangle$ (abbreviated as “Exact” in the legend). For a Gaussian momentum distribution with a finite momentum width $\sigma_{p}=0.05\hbar k_{L}$ and centered around $p_{0}=0$, the BS efficiency is evaluated to be $\eta_{BS}=97.348\%$.
The conventional double Bragg mirror pulse is a temporal Gaussian with parameters $(\Omega_{BS},\tau_{BS},t_{0})=(2.89\,\omega_{\mathrm{rec}},0.64\,\omega_{\mathrm{rec}}^{-1},\,0)$, and a detuning $\Delta=0$. The 2D landscape of the conventional double Bragg mirror efficiency versus quasi-momentum and polarization error, i.e., $\mathcal{F}_{BS}(p,\varepsilon_{pol})$, is shown in Fig. S2(a).
Figure S2: Performance of the C-DBD mirror pulse. (a) Landscape of the C-DBD mirror pulse efficiency $\mathcal{F}_{M}$ versus quasi-momentum $p$ and polarization error $\varepsilon_{pol}$. (b) Transition probabilities to different momentum states after the C-DBD mirror pulse with an input state $|p+2\hbar k_{L}\rangle$ and no polarization error $\varepsilon_{pol}=0$.
In Fig. S2(b), we plot transition probabilities to different momentum states with an input momentum state $|p+2\hbar k_{L}\rangle$ versus quasi-momentum $p$ for a fixed polarization error $\varepsilon_{pol}=0$, and compare it with the exact numerical solutions in position space. For a Gaussian momentum distribution with a finite momentum width $\sigma_{p}=0.05\hbar k_{L}$ and centered around $p_{0}=2\hbar k_{L}$, the mirror efficiency is evaluated to be $\eta_{M}=96.426\%$.
S2.2 CD-DBD beam-splitter pulse
The constant-detuning mitigated DBD (CD-DBD) BS pulse is a temporal Gaussian pulse sharing the same Gaussian parameters as the C-DBD BS pulse, but with an optimized constant detuning $\Delta=0.27\omega_{\mathrm{rec}}$ for quasi-momentum $p=0$ found by a numerical scan. The 2D landscape of the CD-DBD beam-splitter efficiency versus quasi-momentum and polarization error, i.e., $\mathcal{F}_{BS}(p,\varepsilon_{pol})$, is shown in Fig. S3(a).
Figure S3: Performance of the CD-DBD BS pulse. (a) Landscape of the CD-DBD BS pulse efficiency $\mathcal{F}_{BS}$ versus quasi-momentum $p$ and polarization error $\varepsilon_{pol}$. (b) Transition probabilities to different momentum states after the CD-DBD BS pulse with an input state $|p\rangle$ and no polarization error $\varepsilon_{pol}=0$.
In Fig. S3(b), we plot transition probabilities to different momentum states with an input momentum state $|p\rangle$ for a fixed polarization error $\varepsilon_{pol}=0$, and compare it with the exact numerical solutions. For a Gaussian momentum distribution with a finite momentum width $\sigma_{p}=0.05\hbar k_{L}$ and centered around $p_{0}=0$, the BS efficiency is evaluated to be $\eta_{BS}=99.757\%$.
S2.3 DS-DBD beam-splitter and mirror pulses
The linear-detuning-sweep double Bragg (DS-DBD) BS pulse is a temporal Gaussian pulse sharing the same Gaussian parameters as the CD-DBD and C-DBD BS pulse, but with an optimized linear detuning sweep $\Delta(t)/\omega_{\mathrm{rec}}=(0.37/\tau_{BS})(t-t_{0})+0.315$ found by a numerical scan. The 2-dimensional (2D) landscape of the linear-detuning-sweep double Bragg beam-splitter efficiency versus quasi-momentum and polarization error, i.e., $\mathcal{F}_{BS}(p,\varepsilon_{pol})$, is shown in Fig. S4(a).
Figure S4: Performance of the DS-DBD beam-splitter pulse. (a) Landscape of the DS-DBD BS pulse efficiency $\mathcal{F}_{BS}$ versus quasi-momentum $p$ and polarization error $\varepsilon_{pol}$. (b) Transition probabilities to different momentum states after the DS-DBD BS pulse with an input state $|p\rangle$ and no polarization error $\varepsilon_{pol}=0$.
In Fig. S4(b), we plot transition probabilities to different momentum states with an input momentum eigenstate $|p\rangle$ versus quasi-momentum for polarization error $\varepsilon_{pol}=0$, and compare it with the exact numerical solutions. For a Gaussian momentum distribution with a finite momentum width $\sigma_{p}=0.05\hbar k_{L}$ and centered around $p_{0}=0$, the DS-DBD BS efficiency is evaluated to be $\eta_{BS}=99.937\%$.
The DS-DBD mirror pulse is a temporal Gaussian with same parameters $(\Omega_{BS},\tau_{BS},t_{0})=(2.89\,\omega_{\mathrm{rec}},0.64\,\omega_{\mathrm{rec}}^{-1},\,0)$ as the C-DBD M pulse, and a linear detuning $\Delta(t)/\omega_{\mathrm{rec}}=(0.75/\tau_{BS})(t-t_{0})-4$ found by a numerical scan. The 2-dimensional (2D) landscape of the DS-DBD beam-mirror efficiency versus quasi-momentum and polarization error, i.e., $\mathcal{F}_{M}(p,\varepsilon_{pol})$, is shown in Fig. S5(a).
Figure S5: Performance of the DS-DBD mirror pulse. (a) Landscape of the DS-DBD mirror pulse efficiency $\mathcal{F}_{M}$ versus quasi-momentum $p$ and polarization error $\varepsilon_{pol}$. (b) Transition probabilities to different momentum states after the DS-DBD mirror pulse with an input state $|p+2\hbar k_{L}\rangle$ and no polarization error $\varepsilon_{pol}=0$.
In Fig. S5(b), we plot transition probabilities to different momentum states with an input momentum state $|p+2\hbar k_{L}\rangle$ versus quasi-momentum $p$ for a fixed polarization error $\varepsilon_{pol}=0$, and compare it with the exact numerical solutions in position space. For a Gaussian momentum distribution with a finite momentum width $\sigma_{p}=0.05\hbar k_{L}$ and centered around $p_{0}=2\hbar k_{L}$, the DS-mirror efficiency is evaluated to be $\eta_{M}=97.465\%$.
S2.4 OCT-mirror pulse
We want to optimize the robustness against momentum distribution and an initial COM momentum $p_{0}$. To that goal, we allow as optimization parameters: the peak Rabi frequency of the Gaussian envelope, its width, the position of the center of the Gaussian and a smooth time dependent detuning $(\Omega_{M},\tau_{M},t_{0},\Delta(t))$.
We optimize the population transfer of a Mirror interaction by optimizing the average over a range of initial momentum $p\in P=[-0.2\,\hbar k_{L},0.2\,\hbar k_{L}]$, i.e.,
$\mathrm{Cost}_{p}=$ $\bigg{\langle}\Big{|}1-|\langle+2\hbar k_{L}+p|\psi_{-2}(t=t_{f})\rangle_{p}|^{2}\Big{|}$
$+\Big{|}1-|\langle-2\hbar k_{L}+p|\psi_{2}(t=t_{f})\rangle_{p}|^{2}\Big{|}\bigg{\rangle}_{P},$ (S15)
where $|\psi_{\pm 2}(t=t_{f})_{p}\rangle$ stands for the final state after the interaction when the initial state was $|\psi_{\pm 2}(t=0)\rangle_{p}=|\pm 2\hbar k_{L}+p\rangle$.
The OCT-mirror pulse is a temporal Gaussian characterized by the parameters $(\Omega_{M},\tau_{M},t_{0})=(2.502\,\omega_{\mathrm{rec}},1.829\,\omega_{\mathrm{rec}}^{-1},3.879\,\omega_{\mathrm{rec}}^{-1})$, with a time-dependent detuning shown in the upper panel of Fig. S6. The Gaussian envelope is truncated at $T_{\mathrm{pulse}}=10\,\tau_{M}$, and the time axis is shifted as $\tilde{t}=t+T_{\mathrm{pulse}}/2-t_{0}$. The 2D landscape of the OCT-mirror efficiency versus quasi-momentum and polarization error, i.e., $\mathcal{F}_{M}(p,\varepsilon_{pol})$, is shown in the lower panel of Fig. S6.
Figure S6: Smooth OCT-mirror pulse detuning (upper) and the 2D landscape of the OCT-mirror pulse efficiency $\mathcal{F}_{M}$ versus quasi-momentum $p$ and polarization error $\varepsilon_{pol}$ (lower).
Figure S7: Transition probabilities to different momentum states after the OCT-mirror pulse with an input state $|p+2\hbar k_{L}\rangle$ and no polarization error $\varepsilon_{pol}=0$.
In Fig. S7, we plot transition probabilities to different momentum states with an input momentum state $|p+2\hbar k_{L}\rangle$ versus quasi-momentum $p$ for a fixed polarization error $\varepsilon_{pol}=0$, and compare it with the exact numerical solutions in position space. For a Gaussian momentum distribution with a finite momentum width $\sigma_{p}=0.05\hbar k_{L}$ and centered around $p_{0}=2\hbar k_{L}$, the OCT-mirror efficiency is evaluated to be $\eta_{M}=99.806\%$.
We have plotted the time-axis of the detuning in units of $\omega_{\mathrm{rec}}$. To get an idea of the duration in the laboratory, we give an example for the pulse shown in Fig. S6. For the case of $^{87}\text{Rb}$ with a wavelength of $780.1\times 10^{-9}\,m$, the duration is about $760\,\mu s$.

### v2

@@ II.2.4 Mach-Zehnder protocol with optimal control theory
The final detuning-controlled Mach-Zehnder strategy targets the optimization of the mirror pulse—identified as the performance bottleneck in the DS-DBD protocol—by applying optimal control theory (OCT), implemented with QCTRL’s Boulder Opal package Ball et al. (2021). Unlike SBD mirror pulse, the DBD mirror pulse involves a four-photon transition $|\pm 2\hbar k_{L}\rangle\rightarrow|0\hbar k_{L}\rangle\rightarrow|\mp 2\hbar k_{L}\rangle$ making it more sensitive to finite momentum spread and nonzero COM momentum Szigeti et al. (2012); Müller et al. (2008); Chiow et al. (2011); Stolzenberg et al. (2025). As illustrated in Fig. 1, a C-DBD mirror selectively filters and only reflects the central momentum component. This sensitivity is mitigated by jointly optimizing a smooth time-dependent detuning $\Delta(t)$ and all three Gaussian pulse parameters $(\Omega_{M},\,\tau_{M},\,t_{0})$, yielding an OCT mirror pulse with $(\Omega_{M},\,\tau_{M},\,t_{0})=(2.502\,\omega_{\mathrm{rec}},\,1.829\,\omega_{\mathrm{rec}}^{-1},\,3.879\,\omega_{\mathrm{rec}}^{-1})$ and an efficiency of $\eta_{M}=99.806\%$ (see Fig. S8 in Supplementary Material for the OCT detuning $\Delta(t)$), representing a substantial improvement over the DS mirror pulse (see Fig. 3). The resulting hybrid “OCT” sequence combines this mirror with two DS beam-splitters, separated by interrogation time $T$.
@@ III Contrast of double Bragg Mach-Zehnder interferometers
For non-ideal beam-splitter or mirror operations, the contrast is generally less than unity, and the output signal typically contains multiple Fourier components, e.g., Ramsey–Bordé–type parasitic paths produce distinct oscillatory components from those of the main DBD interferometer Kirsten-Siemß et al. (2023). However, the contrast can still be defined, analogous to the ideal case, as the population difference between the first two non-trivial extrema of the output signal when scanning the interrogation time $T$:
@@
where $T_{max}$ and $T_{min}$ denote the first non-trivial maximum and minimum of $P_{\pm 2\hbar k}(g,T)\equiv P_{2}(g,T)+P_{3}(g,T)$ at a fixed $g$ as illustrated by the contrast extraction in Fig. 5(b–c). For instance, Fig. 5(b) shows that the contrast extracted for the DS-DBD strategy for an initial input state with a momentum width of $\sigma_{p}=0.05\hbar k_{L}$ and COM momentum $p_{0}=0$ without polarization error is $\mathbf{C}=97\%$. Fig. 5(c) shows the detrimental effect of a large initial COM momentum, $p_{0}=0.1\hbar k_{L}$, on the contrast, where the contrast is reduced to $\mathbf{C}=83\%$ for a narrower momentum width of $\sigma_{p}=0.01\hbar k_{L}$ without polarization error. In both cases, the results of the five-level S-matrix theory agrees perfectly with the exact numerical solutions of the Schrödinger equation based on the second-order Suzuki-Trotter decomposition Suzuki (1990); Fitzek et al. (2020); Seckmeyer et al. (2025) with a momentum truncation up to $\pm 14.5\hbar k_{L}$. The single Fourier fringe model (Eq. (9)) applies only in the ideal regime of small initial COM momentum and momentum width, where pulse efficiency remains nearly unaffected.
Main results. We compare the contrast robustness of different detuning-control strategies with respect to three dominant experimental imperfections: the atomic momentum width $\sigma_{p}$ [Fig. 5(d)], the initial center-of-mass momentum $p_{0}$ [Fig. 5(e)], and the polarization error $\varepsilon_{pol}$ [Fig. 5(f)]. For realistic experimental conditions with well-controlled polarization errors ($\varepsilon_{pol}<3\%$), we consistently observe the performance hierarchy
$\text{OCT}>\text{DS-DBD}>\text{C-DBD}\approx\text{CD-DBD}.$
@@
The relative improvements of the DS-DBD and OCT protocols over C-DBD (or CD-DBD) depend on the atomic momentum width (set by the 1D effective temperature), the COM momentum (influenced by laser alignment and pulse timing), and the polarization properties of the optical beams. For a typical atom-chip-generated BEC with a 1D effective temperature of $2\,\text{nK}$ Deppner et al. (2021), corresponding to a momentum width of approximately $\sigma_{p}=0.10\hbar k_{L}$ (for $^{87}\text{Rb}$ atoms at a wavelength of $780\,\text{nm}$), the DS-DBD and OCT strategies yield relative contrast improvements of $3.4\%$ and $12.0\%$, respectively, over the C-DBD protocol (assuming perfect polarizations and zero COM momentum). For a BEC after evaporative cooling with a 1D temperature below $500\,\text{pK}$ Leanhardt et al. (2003), or a momentum width about $0.05\hbar k_{L}$, the improvements reduce to $1.3\%$ (for DS-DBD) and $3.8\%$ (for OCT). In the state-of-the-art case of delta-kick-collimated BECs with an effective temperature below $40\,\text{pK}$ Deppner et al. (2021); Kovachy et al. (2015), corresponding to $\sigma_{p}=0.014\hbar k_{L}$, the improvements are $0.4\%$ and $0.5\%$, respectively. With well-controlled polarization error and initial COM momentum, the DS-DBD strategy maintains contrast above $90\%$ up to $\sigma_{p}\leq 0.097\,\hbar k_{L}$, while the OCT strategy remains above $95\%$ contrast up to $\sigma_{p}\leq 0.132\,\hbar k_{L}$.
@@
Figure 6: Contrast as a function of relative peak lattice depth fluctuations for different detuning-control strategies. Symbols with error bars (representing standard deviation) stand for exact numerical simulations; shaded regions denote theoretical confidence intervals from the five-level S-matrix model.
@@ IV Conclusion
In summary, we have proposed and analyzed high-contrast Mach–Zehnder atom interferometers based on double Bragg diffraction (DBD) operated under external acceleration. By introducing a tri-frequency laser configuration that compensates differential Doppler shifts, we enable efficient and symmetric beam-splitter and mirror operations even in non-degenerate geometries. To mitigate contrast loss from realistic experimental imperfections—including momentum spread, center-of-mass motion, polarization errors, and lattice-depth fluctuations—we systematically compared four detuning-control strategies: C-DBD, CD-DBD, DS-DBD, and OCT-DBD. Our analysis demonstrates that properly engineered detuning-control strategies can elevate double Bragg diffraction interferometers to performance levels previously accessible only with Raman-based schemes. In particular, the OCT strategy simultaneously mitigates Doppler shifts, AC-Stark effects, and polarization errors, enabling robust beam-splitter and mirror operations with near-unity efficiency. Results obtained from the five-level S-matrix theory, validated by exact numerical simulations, show that contrasts exceeding $95\%$ can be maintained across realistic momentum widths and laser-power fluctuations—effectively removing a long-standing limitation of DBD interferometry. These findings establish an experimentally feasible pathway toward high-contrast, large-momentum-transfer atom interferometers. By combining these advanced detuning-controlled DBD pulses with subsequent Bloch oscillations—using the experimentally demonstrated BO efficiency of $99.93\%$ per recoil from Ref. Gebbe et al. (2021)—our approach is estimated to yield an overall contrast of roughly $56\%$ for a $408\,\hbar k_{L}$ LMT interferometer. This represents a substantial contrast improvement over state-of-the-art DBD implementations and highlights the potential of detuning-controlled DBD for future high-precision quantum sensors, with applications ranging from terrestrial gravimetry to space-borne tests of fundamental physics.
@@
The online Supplementary Material contains three sections. Sec. S1 provides a detailed derivation of tri-frequency DBD Hamiltonian (Eq. (1)) in the laboratory frame. Sec. S2 provides a detailed derivation of the unitary transformation from the laboratory-frame DBD Hamiltonian (Eq. (1)) to the twin-lattice center-of-mass frame (Eq. (3)). Sec. S3 defines the beam-splitter and mirror pulse efficiencies, evaluates their robustness under experimental imperfections, and outlines the cost function used in the OCT optimization.
@@ acknowledgement
Conflict of interest
The authors have no conflicts to disclose.
Data availability
The data that supports the findings of this study are available from the corresponding author upon reasonable request.